# Recovering contact forms from boundary data

### v5

@@ 4. Contact Holography: recovering contact geometry from the boundary data
Recall that $\beta$ uniquely defines the Reeb field $v_{\beta}$, and the conformal class of $v_{\beta}$ determines the $v_{\beta}$-generated Morse stratification of the boundary.
If $\Phi^{\ast}(\beta)=\beta_{1}$ and $v_{\beta_{1}}=\lambda\cdot v_{\beta}$ for some smooth function $\lambda:X\to\mathbb{R}_{>0}$ (i.e., $\beta_{1},\beta$ are $\mathcal{RC}$-equivalent), then the two Morse stratifications coincide, since they are defined in terms of applying iteratively the Lie derivatives $\mathcal{L}_{v_{\beta}}$ and $\mathcal{L}_{v_{\beta_{1}}}=\lambda\cdot\mathcal{L}_{v_{\beta}}$ to an auxiliary function $h:\hat{X}\to\mathbb{R}$ that determines the boundary $\partial_{1}X$ [K1]. By definition, such a function $h$ has the three properties:
[…]
Lemma 4.7 states a relative version of the Gray’s theorem [Gray]. Its claim may be found in [EM], page 95. Lemma 5.5 will generalize a bit Lemma 4.7 (whose proof we skip).
Lemma 4.7.
@@
Corollary 4.7.
Let contact forms $\beta_{1}$ and $\beta_{2}$ on $X$ generate contact structures $\xi_{1}$ and $\xi_{2}$ which belong to the same path-connected component of the space $\mathsf{Cont}(X,Op(\xi^{\partial}))$. Assume that the Reeb vector field $v_{1}$ of the contact form $\beta_{1}$, is traversing.
Then the Reeb vector field $v_{2}$ of a contact form $\beta_{2}$, is traversing as well. As a result, any $v_{2}$-trajectory is homeomorphic to a closed interval or to a singleton (in particular, $v_{2}$ has no closed or trapped trajectories).
Proof.
Consider a smooth path in $\mathsf{Cont}(X,Op(\xi^{\partial}))$ that connects $\xi_{1}$ to $\xi_{2}$. By Lemma 4.7, there exists a smooth orientation-preserving diffeomorphism $\Phi$, identical on $Op(\partial_{1}X)$, such that $\lambda\cdot\beta_{1}=\Phi^{\ast}(\beta_{2})$ for a smooth positive function $\lambda:X\to\mathbb{R}_{+}$. Let us check that $v_{2}=_{\mathsf{def}}\Phi_{\ast}(\lambda^{-1}v_{1})$ is the Reeb vector field for $\beta_{2}$:
$v_{2}\,\rfloor\;d\beta_{2}=\Phi_{\ast}(\lambda^{-1}v_{1})\,\rfloor\;d\beta_{2}=\lambda^{-1}v_{1}\,\rfloor\;\Phi^{\ast}(d\beta_{2})=v_{1}\,\rfloor\;\Big{(}\frac{d\lambda}{\lambda}\wedge\beta_{1}+d\beta_{1}\Big{)}=v_{1}\,\rfloor\;\Big{(}\frac{d\lambda}{\lambda}\wedge\beta_{1}\Big{)}.$
Since $\big{(}\frac{d\lambda}{\lambda}\wedge\beta_{1}\big{)}\big{|}_{\ker(\beta_{1})}=0$, the $1$-form $(v_{2}\,\rfloor\,d\beta_{2})|_{\ker\beta_{1}}=0$. Also, $(v_{2}\,\rfloor\,d\beta_{2})$ vanishes on $v_{1}$:
$v_{1}\,\rfloor(v_{2}\,\rfloor\,d\beta_{2})=v_{1}\,\rfloor\,\big{(}v_{1}\,\rfloor\,\big{(}\frac{d\lambda}{\lambda}\wedge\beta_{1}\big{)}\big{)}=0.$
Therefore, $v_{2}\,\rfloor\,d\beta_{2}=0$. Moreover,
$\beta_{2}(v_{2})=\beta_{2}(\Phi_{\ast}(\lambda^{-1}\cdot v_{1}))=\Phi^{\ast}(\beta_{2})(\lambda^{-1}\cdot v_{1})=\lambda\cdot\beta_{1}(\lambda^{-1}\cdot v_{1})=1.$
Hence, $v_{2}=\Phi_{\ast}(\lambda\cdot v_{1})$ is the Reeb vector field for the form $\beta_{2}$. Therefore, the Reeb vector field $v_{2}$ is traversing as well. ∎
@@
Let $X_{1},X_{2}$ be compact connected oriented smooth $(2n+1)$-dimensional manifolds with boundaries, equipped with contact forms $\beta_{1},\beta_{2}$. We assume that their Reeb vector fields $v_{1},v_{2}$ are traversing and boundary generic. We denote by $f_{1}^{\bullet},f_{2}^{\bullet}$ the Lyapunov functions such that $df_{1}^{\bullet}(v_{1})=1=df_{2}^{\bullet}(v_{2})$, whose existence is guarantied by Lemma 4.1. In addition, we assume that the $v_{2}$ possesses Property $\mathsf{A}$ from Definition 3.1.222Recall that the property $\partial_{3}X_{2}(v_{2})=\emptyset$ implies Property $\mathsf{A}$.
Let a smooth orientation-preserving diffeomorphism $\Phi^{\partial}:\partial_{1}X_{1}\to\partial_{1}X_{2}$ commute with the two causality maps:
@@
$\bullet$ Then $\Phi^{\partial}$ extends to a smooth orientation-preserving diffeomorphism $\Phi:X_{1}\to X_{2}$ such that $\Phi$ maps the oriented foliation $\mathcal{F}(v_{1})$ to the oriented foliation $\mathcal{F}(v_{2})$, and the pull-back contact form $\beta^{\dagger}_{1}=\Phi^{\ast}(\beta_{2})$ is $\mathcal{RC}$-equivalent to $\beta_{1}$ (in the sense of Definition 4.5).
@@
Conversely, using the traversing nature of $v_{i}$, any pair $(y,\,t)$, where $y\in\gamma_{x}\cap\partial_{1}X_{i}$ and $t\in[f_{i}^{\partial}(\gamma_{x}\cap\partial_{1}X_{i})]\subset\mathbb{R}$ determines a unique point $x\in X_{i}$. Here $[f_{i}^{\partial}(\gamma_{x}\cap\partial_{1}X_{i})]$ denotes the minimal closed interval that contains the finite set $f_{i}^{\partial}(\gamma_{x}\cap\partial_{1}X_{i})$. Note that some pairs of leaves $L$ and $\gamma$ may have an empty intersection, and some components of leaves $L$ may have an empty intersection with the boundary $\partial_{1}X_{i}$. However, if $f_{i}(\partial_{1}X_{i})=f_{i}(X_{i})$ as sets, then $f_{i}^{-1}(c)\cap\partial_{1}X_{i}\neq\emptyset$ for any $c\in f_{i}(X_{i})$. In particular, if $\partial_{1}X_{i}$ is connected, then $f_{i}(\partial_{1}X_{i})=f_{i}(X_{i})$ and $f_{i}^{-1}(c)\cap\partial_{1}X_{i}\neq\emptyset$ for any $c\in f_{i}(X_{i})$
@@
By the arguments as in the proof of Corollary 4.7, $v_{1}^{\dagger}=\Phi^{-1}_{\ast}(v_{2})$ is the Reeb vector field for $\beta_{1}^{\dagger}$. Thus, $\beta_{1}^{\dagger}$ is $v_{1}^{\dagger}$-invariant. Since $v_{1}^{\dagger}=\lambda\cdot v_{1}$, where $\lambda:X_{1}\to\mathbb{R}_{+}$ is a positive smooth function, $\beta_{1}$ and $\beta_{1}^{\dagger}$ are $\mathcal{RC}$-equivalent contact forms, and $\xi_{1}$ and $\xi_{1}^{\dagger}$ are $\mathcal{RC}$-equivalent contact structures on $X_{1}$ (see Definition 4.5). Equivalently, $\xi_{2}^{\dagger}=_{\mathsf{def}}\Phi_{\ast}(\xi_{1})$ is $\mathcal{RC}$-equivalent to $\xi_{2}$. This proves the first assertion of the theorem.
Now we turn to the second assertion. The special Lypunov functions $f_{1}^{\bullet},f_{2}^{\bullet}$ are essential for constructing the desired extension $\Phi$ of $\Phi^{\partial}$. We can build the extension $\Phi$ of $\Phi^{\partial}$ as before, using the map $\Phi^{\mathcal{T}}:\mathcal{T}(v_{1})\to\mathcal{T}(v_{2})$ (induced by $\Phi^{\partial}$) and the restrictions $(f_{1}^{\bullet})^{\partial}=(\Phi^{\partial})^{\ast}((f_{2}^{\bullet})^{\partial})$ and $(f_{2}^{\bullet})^{\partial}$ of the Lyapunov functions $f_{1}^{\bullet}$ and $f_{2}^{\bullet}$. By the very construction of $\Phi$, we get $\Phi^{\ast}(f_{2}^{\bullet})=f_{1}^{\bullet}$. Since $df_{i}^{\bullet}(v_{i})=1$ and $\Phi$ maps $v_{1}$-trajectories to $v_{2}$-trajectories, it follows that $\Phi_{\ast}(v_{1})=v_{2}$. By the hypotheses, $(\Phi^{\partial})^{\ast}(\beta_{2}|_{\partial_{1}X_{2}})=\beta_{1}|_{\partial_{1}X_{1}}$ and, since $\beta_{1}(v_{1})=1=\beta_{2}(v_{2})$, we conclude that $(\Phi^{\partial})^{\ast}(\beta_{2}||_{\partial_{1}X_{2}})=\beta_{1}||_{\partial_{1}X_{1}}$, except, perhaps, along the locus $\partial_{2}X_{2}$ where $v_{2}$ is tangent to $\partial_{1}X_{2}$. However, since $\Phi^{\partial}$ is a smooth diffeomorphism, by continuity, we get that $(\Phi^{\partial})^{\ast}(\beta_{2}||_{\partial_{1}X_{2}})=\beta_{1}||_{\partial_{1}X_{1}}$ without exception. Thus, using that $\beta_{i}$ is $v_{i}$-invariant, we get that $(\Phi)^{\ast}(\beta_{2})=\beta_{1}$. ∎
Note that this construction of the extension $\Phi$ from the first assertion of Theorem 4.1 is implicit and far from being canonical. For example, it depends on the choice of extension of $f_{1}^{\partial}=_{\mathsf{def}}(\Phi^{\partial})^{\ast}(f_{2}^{\partial})$ to a smooth function $f_{1}:X_{1}\to\mathbb{R}$, which is strictly monotone along the $v_{1}$-trajectories. The uniqueness (rigidity) of the extension $\Phi$ may be achieved, if we assume the full knowleadge of the manifolds $X_{i}$, equipped with the foliation grids $\mathcal{F}(v_{i}),\mathcal{G}(f_{i})$ and the Lyapunov functions $f_{i}$ or $f_{i}^{\bullet}$ (see Corollary 4.8).
@@
Let $X$ be a compact connected smooth $(2n+1)$-dimensional manifold with boundary, equipped with a contact form $\beta$. We assume that the Reeb vector field $v=v_{\beta}$ is traversing and boundary generic and possesses Property $\mathsf{A}$. By Lemma 4.1, using $\beta$, we construct a Lyapunov function $f^{\bullet}:X\to\mathbb{R}$ such that $df^{\bullet}(v)=1$.
@@
are sufficient for a reconstruction of the triple $(X,f^{\bullet},\beta)$ (and thus, of the Reeb vector field $v_{\beta}$ as well), up to diffeomorphisms $\Phi:X\to X$ which are the identity map on $\partial_{1}X$.
@@
Assume that we have two sets of structures, $\{X,\,\mathcal{F}(v_{\beta_{1}}),\,\mathcal{G}(f^{\bullet}_{1}),\beta_{1}\}$ and $\{X,\,\mathcal{F}(v_{\beta_{2}}),\,\hfill\break\mathcal{G}(f^{\bullet}_{2}),\beta_{2}\}$, which agree along $\partial_{1}X$. By Theorem 4.1, being applied to $\Phi^{\partial}=\mathsf{id_{\partial_{1}X}}$, the following data
@@
Although the construction of an extension $\Phi:X\to X$ of a diffeomorphism $\Phi^{\partial}:\partial_{1}X\to\partial_{1}X$, subject to the hypotheses of Corollary 4.8, is not canonic, we will argue that by fixing one such extension $\Phi_{\star}$, makes it possible to extend any other diffeomorphism $\tilde{\Phi}^{\partial}:\partial_{1}X\to\partial_{1}X$, subject to the hypotheses of Corollary 4.8, in a rigid way, so that the reconstruction works well for families of such $\tilde{\Phi}^{\partial}$’s.
We assume now that the quadruple $(X,\mathcal{F}(v_{\beta}),f^{\bullet},\beta)$ is known and fixed. We follow the scheme in (3). Then, given any diffeomorphism $\Phi^{\partial}:\partial_{1}X\to\partial_{1}X$ that respects the boundary data (4.4), we can extend it canonically to a diffeomorphism $\Phi:X\to X$ that keeps $\{\mathcal{F}(v),f^{\bullet},\beta\}$ invariant. Indeed, any point $x\in X$ is determined by the unique pair $\big{(}\gamma_{x}\cap\partial_{1}X,\;f^{\bullet}(x)\big{)}$, subject to the requirement $f^{\bullet}(x)\in[f^{\bullet}(\gamma_{x}\cap\partial_{1}X)]$. Now put $\Phi(x)=x^{\prime}$, where the point $x^{\prime}\in X$ is uniquely defined by the pair
@@
Assuming the validity of Property $\mathsf{A}$, as in Theorem 4.1, we get that $\Phi$ is a smooth diffeomorphism.
@@
Consider the topological groups $\mathsf{Diff}^{\infty}_{+}(X)$ and $\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$ of smooth orientation-preserving diffeomorphisms and the natural restriction homomorphism $\mathcal{R}:\mathsf{Diff}^{\infty}_{+}(X)\to\mathsf{Diff}^{\infty}_{+}(\partial_{1}X).$ Similarly, we may consider the groups ${}^{\circ}\mathsf{Diff}^{\infty}_{+}(X)$ and ${}^{\circ}\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$ that are the connected components of $\mathsf{Diff}^{\infty}_{+}(X)$ and $\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$ that contain the identity.
Recall that the famous Cerf Theorem [Cerf] claims that any orientation-preserving diffeomorphism of $S^{3}$ is isotopic to the identity, and thus, extends to a diffeomorphism of the $4$-ball $D^{4}$. Therefore, the homomorphism $\mathcal{R}:\mathsf{Diff}^{\infty}_{+}(D^{4})\to\mathsf{Diff}^{\infty}_{+}(S^{3})$ is onto.
However, remarkably, for any smooth compact connected $X$ with boundary, the restriction map $\mathcal{R}:\;^{\circ}\mathsf{Diff}^{\infty}_{+}(X)\to\;^{\circ}\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$ does not admit a group-theoretical section ([ChMa], Corollary 1.8)!
@@
Let $\mathsf{Diff}^{\infty}_{+}(\partial_{1}X;\,C_{v_{\beta}},\beta^{\partial},(f^{\bullet})^{\partial})$ denote the subgroup of $\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$ whose elements commute with the causality map $C_{v_{\beta}}:\partial_{1}^{+}X(v_{\beta})\to\partial_{1}^{-}X(v_{\beta})$, preserve the restriction to the boundary $(f^{\bullet})^{\partial}:\partial_{1}X\to\mathbb{R}$ of the Lyapunov function $f^{\bullet}$ and the section $\beta^{\partial}:\partial_{1}X\to T^{\ast}(X)|_{\partial_{1}X}$.

### v6

@@ 4. Contact Holography: recovering contact geometry from the boundary data
Recall that $\beta$ uniquely defines the Reeb field $v_{\beta}$, and the conformal class of $v_{\beta}$ determines the $v_{\beta}$-generated Morse stratification of the boundary [K1].
If $\Phi^{\ast}(\beta)=\beta_{1}$ and $v_{\beta_{1}}=\lambda\cdot v_{\beta}$ for some smooth function $\lambda:X\to\mathbb{R}_{>0}$ (i.e., $\beta_{1},\beta$ are $\mathcal{RC}$-equivalent), then the two Morse stratifications coincide, since they are defined in terms of applying iteratively the Lie derivatives $\mathcal{L}_{v_{\beta}}$ and $\mathcal{L}_{v_{\beta_{1}}}=\lambda\cdot\mathcal{L}_{v_{\beta}}$ to the same auxiliary function $h:\hat{X}\to\mathbb{R}$ that determines the boundary $\partial_{1}X$ [K1]. By definition, such a function $h$ has the three properties:
$\bullet\;$ $0\text{ is a regular value of }h,$
$\bullet\;$ $h^{-1}(0)=\partial_{1}X$
(4.3) $\bullet\;$ $h^{-1}((-\infty,0])=X.$
For $\lambda>0$, by an inductive argument in $j$, the loci
$\partial_{j}X(v_{\beta})=\big\{x\in\partial X\big|\;h(x)=0,(\mathcal{L}_{v_{\beta}}h)(x)=0,\ldots,(\mathcal{L}^{j-1}_{v_{\beta}}h)(x)=0\big\}$
$\partial_{j}X(\lambda\cdot v_{\beta})=\big\{x\in\partial X\big|\;h(x)=0,(\mathcal{L}_{\lambda\cdot v_{\beta}}h)(x)=0,\ldots,(\mathcal{L}^{j-1}_{\lambda\cdot v_{\beta}}h)(x)=0\big\}$
coincide. ∎
@@
Let $X_{1},X_{2}$ be compact connected oriented smooth $(2n+1)$-dimensional manifolds with boundaries, equipped with contact forms $\beta_{1},\beta_{2}$. We assume that their Reeb vector fields $v_{1}=_{\mathsf{def}}v_{\beta_{1}},\,v_{2}=_{\mathsf{def}}v_{\beta_{2}}$ are traversing and boundary generic. We denote by $f_{1}^{\bullet},f_{2}^{\bullet}$ the Lyapunov functions such that $df_{1}^{\bullet}(v_{1})=1=df_{2}^{\bullet}(v_{2})$ (their existence is guarantied by Lemma 4.1). In addition, we assume that the $v_{2}$ possesses Property $\mathsf{A}$ from Definition 3.1.333Recall that the property $\partial_{3}X_{2}(v_{2})=\emptyset$ implies Property $\mathsf{A}$.
Assume that a smooth orientation-preserving diffeomorphism $\Phi^{\partial}:\partial_{1}X_{1}\to\partial_{1}X_{2}$ commutes with the two causality maps:
@@
$\bullet$ Then $\Phi^{\partial}$ extends to a smooth orientation-preserving diffeomorphism $\Phi:X_{1}\to X_{2}$ such that $\Phi$ maps the oriented foliation $\mathcal{F}(v_{1})$ to the oriented foliation $\mathcal{F}(v_{2})$, and the pull-back contact form $\beta^{\dagger}_{1}=\Phi^{\ast}(\beta_{2})$ is $\mathcal{RC}$-equivalent (in the sense of Definition 4.5) to $\beta_{1}$.
@@
Conversely, using the traversing nature of the vector field $v_{i}$, any pair (y,t)(y,\,t), where $y\in\gamma_{x}\cap\partial_{1}X_{i}$ and $t\in[f_{i}^{\partial}(\gamma_{x}\cap\partial_{1}X_{i})]\subset\mathbb{R}$ determines a unique point $x\in X_{i}$. Here $f_{i}^{\partial}(\gamma_{x}\cap\partial_{1}X_{i})]$ denotes the minimal closed interval that contains the finite set $f_{i}^{\partial}(\gamma_{x}\cap\partial_{1}X_{i})$. Note that some pairs of leaves $L$ and $\gamma$ may have an empty intersection, and some components of leaves $L$ may have an empty intersection with the boundary $\partial_{1}X_{i}$. However, if $f_{i}(\partial_{1}X_{i})=f_{i}(X_{i})$ as sets, then $f_{i}^{-1}(c)\cap\partial_{1}X_{i}\neq\emptyset$ for any $c\in f_{i}(X_{i})$. In particular, if $\partial_{1}X_{i}$ is connected, then $f_{i}(\partial_{1}X_{i})=f_{i}(X_{i})$ and $f_{i}^{-1}(c)\cap\partial_{1}X_{i}\neq\emptyset$ for any $c\in f_{i}(X_{i})$
@@
In fact, $v_{1}^{\dagger}=\Phi^{-1}_{\ast}(v_{2})$ is the Reeb vector field for $\beta_{1}^{\dagger}$. Indeed, by naturality, $\Phi^{\ast}(\beta_{2})(v_{1}^{\dagger})=\beta_{2}(v_{2})=1$ and $\Phi^{\ast}(d\beta_{2})(v_{1}^{\dagger}\wedge\sim)=d\beta_{2}(v_{2}\wedge\sim)=0$. Thus, $\beta_{1}^{\dagger}$ is $v_{1}^{\dagger}$-invariant. Since $v_{1}^{\dagger}=\lambda\cdot v_{1}$, where $\lambda:X_{1}\to\mathbb{R}_{+}$ is a positive smooth function, $\beta_{1}$ and $\beta_{1}^{\dagger}$ are $\mathcal{RC}$-equivalent contact forms, and $\xi_{1}$ and $\xi_{1}^{\dagger}$ are $\mathcal{RC}$-equivalent contact structures on $X_{1}$ (see Definition 4.5). Equivalently, $\xi_{2}^{\dagger}=_{\mathsf{def}}\Phi_{\ast}(\xi_{1})$ is $\mathcal{RC}$-equivalent to $\xi_{2}$. This proves the first assertion of the theorem.
Now we turn to the second assertion. The special Lypunov functions $f_{1}^{\bullet},f_{2}^{\bullet}$ are essential for constructing the desired extension $\Phi$ of $\Phi^{\partial}$. We can build the extension $\Phi$ of $\Phi^{\partial}$ as before, using the map $\Phi^{\mathcal{T}}:\mathcal{T}(v_{1})\to\mathcal{T}(v_{2})$ (induced by $\Phi^{\partial}$) and the restrictions $(f_{1}^{\bullet})^{\partial}=(\Phi^{\partial})^{\ast}((f_{2}^{\bullet})^{\partial})$ and $(f_{2}^{\bullet})^{\partial}$ of the Lyapunov functions $f_{1}^{\bullet}$ and $f_{2}^{\bullet}$ to the corresponding boundaries. By the very construction of $\Phi$, we get $\Phi^{\ast}(f_{2}^{\bullet})=f_{1}^{\bullet}$. Since $df_{i}^{\bullet}(v_{i})=1$ and $\Phi$ maps $v_{1}$-trajectories to $v_{2}$-trajectories, it follows that $\Phi_{\ast}(v_{1})=v_{2}$. By the hypotheses, $(\Phi^{\partial})^{\ast}(\beta_{2}|_{\partial_{1}X_{2}})=\beta_{1}|_{\partial_{1}X_{1}}$ and, since $\beta_{1}(v_{1})=1=\beta_{2}(v_{2})$, we conclude that $(\Phi^{\partial})^{\ast}(\beta_{2}||_{\partial_{1}X_{2}})=\beta_{1}||_{\partial_{1}X_{1}}$, except, perhaps, along the locus $\partial_{2}X_{2}$ where $v_{2}$ is tangent to $\partial_{1}X_{2}$. However, since $\Phi^{\partial}$ is a smooth diffeomorphism, by continuity, we get that $(\Phi^{\partial})^{\ast}(\beta_{2}||_{\partial_{1}X_{2}})=\beta_{1}||_{\partial_{1}X_{1}}$ without exception. Thus, using that $\beta_{i}$ is $v_{i}$-invariant, we get that $(\Phi)^{\ast}(\beta_{2})=\beta_{1}$. ∎
Note that this construction of the extension $\Phi$ from the first assertion of Theorem 4.1 is implicit and far from being canonical. For example, it depends on the choice of extension of $f_{1}^{\partial}=_{\mathsf{def}}(\Phi^{\partial})^{\ast}(f_{2}^{\partial})$ to a smooth function $f_{1}:X_{1}\to\mathbb{R}$, which is strictly monotone along the $v_{1}$-trajectories. The uniqueness (rigidity) of the extension $\Phi$ may be achieved, if we assume the full a priori knowleadge of the manifolds $X_{i}$, equipped with the foliation grids $\mathcal{F}(v_{i}),\mathcal{G}(f_{i})$ and the Lyapunov functions $f_{i}$ or $f_{i}^{\bullet}$ (see Corollary 4.8).
@@
Let $X$ be a compact connected smooth $(2n+1)$-dimensional manifold with boundary, equipped with a contact form $\beta$. We assume that the Reeb vector field $v=v_{\beta}$ is traversing, boundary generic, and possesses Property $\mathsf{A}$. By Lemma 4.1, using $\beta$, we construct a Lyapunov function $f^{\bullet}:X\to\mathbb{R}$ such that $df^{\bullet}(v)=1$.
@@
are sufficient for a reconstruction of the triple $(X,f^{\bullet},\beta)$ (and thus, of the Reeb vector field $v_{\beta}$), up to diffeomorphisms $\Phi:X\to X$ which are the identity map on $\partial_{1}X$.
@@
Assume that we have two sets of structures, $\{X,\,\mathcal{F}(v_{\beta_{1}}),\,\mathcal{G}(f^{\bullet}_{1}),\beta_{1}\}$ and $\{X,\,\mathcal{F}(v_{\beta_{2}}),\,\hfill\penalty-10000\mathcal{G}(f^{\bullet}_{2}),\beta_{2}\}$, which agree along $\partial_{1}X$. By Theorem 4.1, being applied to $\Phi^{\partial}=\mathsf{id_{\partial_{1}X}}$, the following data
@@
Although the construction of an extension $\Phi:X\to X$ of a diffeomorphism $\Phi^{\partial}:\partial_{1}X\to\partial_{1}X$, subject to the hypotheses of Corollary 4.8, is not canonical, we will argue that by fixing one such extension $\Phi_{\star}$, makes it possible to extend any other diffeomorphism $\tilde{\Phi}^{\partial}:\partial_{1}X\to\partial_{1}X$, subject to the hypotheses of Corollary 4.8, in a rigid way, so that our reconstruction works well for families of such $\tilde{\Phi}^{\partial}$’s.
We assume now that the quadruple $(X,\mathcal{F}(v_{\beta}),f^{\bullet},\beta)$ is known and fixed. We follow the scheme in (3). Then, given any diffeomorphism $\Phi^{\partial}:\partial_{1}X\to\partial_{1}X$ that respects the boundary data (4.4), we can extend it canonically to a diffeomorphism $\Phi:X\to X$ that keeps the triple $\{\mathcal{F}(v),f^{\bullet},\beta\}$ invariant. Indeed, any point $x\in X$ is determined by the unique pair $\big(\gamma_{x}\cap\partial_{1}X,\;f^{\bullet}(x)\big)$, subject to the requirement $f^{\bullet}(x)\in[f^{\bullet}(\gamma_{x}\cap\partial_{1}X)]$. Now put $\Phi(x)=x^{\prime}$, where the point $x^{\prime}\in X$ is uniquely defined by the pair
@@
Assuming Property $\mathsf{A}$, as in Theorem 4.1, we get that $\Phi$ is a smooth diffeomorphism.
@@
Now, consider the topological groups $\mathsf{Diff}^{\infty}_{+}(X)$ and $\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$ of smooth orientation-preserving diffeomorphisms and the natural restriction homomorphism $\mathcal{R}:\mathsf{Diff}^{\infty}_{+}(X)\to\mathsf{Diff}^{\infty}_{+}(\partial_{1}X).$ Similarly, we may consider the groups $^{\circ}\mathsf{Diff}^{\infty}_{+}(X)$ and $^{\circ}\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$ that are the connected components of $\mathsf{Diff}^{\infty}_{+}(X)$ and $\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$ that contain the identity.
Recall that the famous Cerf Theorem [Cerf] claims that any orientation-preserving diffeomorphism of $S^{3}$ is isotopic to the identity (i.e., $\mathsf{Diff}^{\infty}_{+}(S^{3})\approx\;^{\circ}\mathsf{Diff}^{\infty}_{+}(S^{3})$), and thus, extends to a diffeomorphism of the $4$-ball $D^{4}$. Therefore, the homomorphism $\mathcal{R}:\mathsf{Diff}^{\infty}_{+}(D^{4})\to\mathsf{Diff}^{\infty}_{+}(S^{3})$ is onto.
However, remarkably, for any smooth compact connected manifold $X$ with boundary, the restriction map $\mathcal{R}:\;^{\circ}\mathsf{Diff}^{\infty}_{+}(X)\to\;^{\circ}\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$ does not admit a group-theoretical section ([ChMa], Corollary 1.8)!
@@
Let $\mathsf{Diff}^{\infty}_{+}(\partial_{1}X;\,C_{v_{\beta}},\beta^{\partial},(f^{\bullet})^{\partial})$ denote the subgroup of $\mathsf{Diff}^{\infty}_{+}(\partial_{1}X)$ whose elements commute with the causality map $C_{v_{\beta}}:\partial_{1}^{+}X(v_{\beta})\to\partial_{1}^{-}X(v_{\beta})$, preserve the restrictions to the boundary $(f^{\bullet})^{\partial}:\partial_{1}X\to\mathbb{R}$ of the Lyapunov function $f^{\bullet}$ and the restriction $\beta^{\partial}:\partial_{1}X\to T^{\ast}(X)|_{\partial_{1}X}$ of $\beta:X\to T^{\ast}(X)$.